# Diving into a multi-band holographic superconductor

### v2

@@ 3.4 Thermal a-function:
Moreover, the thermal a-function can be proved to be monotonically decreasing by using null energy condition.
We choose to still use the coordinate in Eq.(7) to express the thermal a-function. In four dimensional case, the thermal a-function simplifies to be
@@
Note that the monotonically decreasing behavior of thermal a-function can also been seen from the monotonically increasing behavior of $\chi$ field. Near the singularity, we found that the thermal a-function becomes
$a_{T}(z)|_{kasner}=\frac{2\pi}{\kappa^{2}}e^{-(\alpha_{1}^{2}+\alpha_{2}^{2})\ln z}=\frac{2\pi}{\kappa^{2}}z^{-\frac{6p_{t}+2}{1-p_{t}}}$ (51)
where we use Eq.(28). Thus we see that the thermal a-function in Kasner region is solely determined by the Kasner exponent $p_{t}$ and different behavior of Kasner exponent will lead to different behavior of thermal a-function (or different behavior of effective degrees of freedom). So the thermal a-function is a good physical quantity to reflect the different interior structure between multi-band superconductor model and single band superconductor model.
4 Conclusion and future direction
In this work, we investigate the interior structure of holographic multi-band superconductor. Besides the boundary phase diagram, the black hole interior structures are also modified by the additional scalar field. There is ER collapse epoch after which Josephson oscillation occurs. When lowering the temperature below which second scalar condensate appears, there will be oscillation behavior of two scalar fields. Moreover, the additional scalar field also has contributions to the Kasner exponent which makes the Kasner exponent different from the single scalar case. The Kasner transition behavior will also be more complicated in double scalar case, which presents more complex relations compared to the "Kasner inversion" phenomenon in single scalar case. Finally, we show that the difference between the Kasner exponent $p_{t}$ can also be reflected in the thermal a-function. These results show that the difference between boundary field theory systems can be reflected in terms of black hole interior structure especially the structure near the black hole singularity.
There are several interesting directions to be pursued in the future investigations. Firstly, we only investigate the multi-band holographic superconductor model for two s-wave order parameters case. As p-wave superconductor has more intricate interior structure compared with s-wave case Cai:2021obq . Thus it is also interesting to investigate the interior of multi-band model with coexistence of s-wave and p-wave order parameters Nie:2013sda . Moreover, besides thermal a-function, there are also proposals using correlation functions Frenkel:2020ysx ; Grinberg:2020fdj ; David:2022nfn and complexity Jorstad:2023kmq to characterize the information of black hole interior (especially the information near the singularity), it is also intriguing to see the behaviors of these variables in holographic multi-band superconductor model.
Acknowledgements.
We are grateful for the useful discussions with our group members. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos.12405066, 12175105 and 11965013. YSA is also supported by the Natural Science Foundation of Jiangsu Province under Grant No. BK20241376.
@@ Appendix A More Kasner transition diagram:
Figure 12: Kasner inversion behavior for $T=0.72292Tc$, $T=0.72275Tc$, $T=0.72265Tc$ and $T=0.72261T_{c}$, all of these example satisfy the two relations in Eq.(46)

### v3

@@ 4 Conclusions and Discussions
Note that the thermal a-function reduces to central charge of boundary CFT for $z\to 0$. Moreover, the monotonicity of thermal a-function can be proved by using null energy condition.
In four dimensional case, the thermal a-function simplifies to be  [^7] Note that the monotonically decreasing behavior of thermal a-function can also been seen from the monotonically increasing behavior of $\chi$ field.
@@
Near the singularity, we found that the thermal a-function becomes
$a_{T}(z)|_{\mathrm{kasner}}=\frac{2\pi}{\kappa^{2}}e^{-(\alpha_{1}^{2}+\alpha_{2}^{2})\ln z}=\frac{2\pi}{\kappa^{2}}z^{-\frac{6p_{t}+2}{1-p_{t}}},$ (39)
where we use Eq.(19). Thus we see that the thermal a-function in Kasner region is solely determined by the Kasner exponent $p_{t}$ and different behavior of Kasner exponent will lead to different behavior of thermal a-function (or different behavior of effective degrees of freedom). So the thermal a-function is a good physical quantity to reflect the different near singularity structure between multi-band superconductor model and single band superconductor model. Besides thermal a-function, there are also proposals using correlation functions [47, 48, 49, 50] and complexity [51] to characterize the near singularity structure of black hole interior, we leave the analysis of these variables to future investigations.
Acknowledgements
We are grateful for the useful discussions with our group members. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos.12405066, 12175105 and 11965013. YSA is also supported by the Natural Science Foundation of Jiangsu Province under Grant No. BK20241376 and Fundamental Research Funds for the Central Universities.
@@ Appendix A More Kasner transition diagram:
Figure 9: Kasner inversion behavior for $T=0.72292Tc$, $T=0.72275Tc$, $T=0.72265Tc$ and $T=0.72261T_{c}$, all of these example satisfy the two relations in Eq.(34).